# Constraining the Luminosity Function and Delay-Time Distribution of Short Gamma-Ray Bursts for Multimessenger Gravitational-Wave Detection Rate Estimation

### v2

@@ 4.1 GW detection rate simulation
We assume that the mass distribution of the NSs obeys a normal distribution with a mean of 1.33 $\mathrm{M_{\odot}}$ and a standard deviation of 0.09 $\mathrm{M_{\odot}}$ (Özel & Freire, 2016). The spins of the NSs are uniformly distributed between $0.050<\chi<$ 0.05, with the upper limit set by the maximum known NS spin (Zhu et al., 2018). The direction of the orbital angular momentum is uniformly distributed in space. We generate the GW waveforms using the Python package PyCBC [^3] https://pycbc.org/ with the IMPRPhenomPv2 approximant model (Schmidt et al., 2012; Hannam et al., 2014; Khan et al., 2016).
@@
[…]
@@
SVOM/$\mathrm{GRM^{(a)}}$ 1–1000 $3.2\,(\mathrm{photons\,cm^{-2}\,s^{-1}})$ 0.24 0.85
EP/$\mathrm{WXT^{(b)}}$ 0.5–4 $8.9\times 10^{-10}\,(\mathrm{erg\,cm^{-2}\,s^{-1}})$ 0.09 0.67
@@
Table 4: The sGRB and joint sGRB $+$ GW detection rates for all combinations of high-energy GRB missions, GW networks, and delay-time distribution models. Columns 2-4 give the sGRB detection rates for the three GRB missions, independent of GW detection. Note that these rates are estimated for events with $d_{L}<1.6\,\mathrm{Gpc}$ (corresponding to $z<0.3$), which is the distance limit of this work; they should not be confused with the total sGRB detection rates (see the discussion at the end of section 4.2). The last six columns provide the joint sGRB $+$ GW detection rates. For instance, the last cell of this table states that for the Gaussian delay model, EP/WXT $+$ IGWN5 will detect $0.24_{-0.04}^{+0.03}$ events per year. All the results are presented at the $1\sigma$ confidence level.
@@
Gaussian $\mathrm{3.28_{-0.58}^{+0.48}}$ $2.69_{-0.40}^{+0.32}$ $0.78^{+0.09}_{-0.11}$ $0.24_{-0.03}^{+0.03}$ $0.09_{-0.01}^{+0.01}$ $0.02_{-0.01}^{+0.01}$ $1.18_{-0.21}^{+0.17}$ $0.69_{-0.12}^{+0.10}$ $0.24_{-0.04}^{+0.03}$
Lognormal $2.82_{-0.31}^{+0.22}$ $2.06_{-0.22}^{+0.16}$ $0.76_{-0.06}^{+0.08}$ $0.21_{-0.02}^{+0.03}$ $0.08_{-0.01}^{+0.01}$ $0.02_{-0.01}^{+0.01}$ $1.03_{-0.10}^{+0.19}$ $0.59_{-0.06}^{+0.10}$ $0.21_{-0.02}^{+0.01}$
Power law $3.19_{-0.35}^{+0.31}$ $1.99_{-0.25}^{+0.28}$ $0.57_{-0.07}^{+0.08}$ $0.21_{-0.02}^{+0.02}$ $0.08_{-0.01}^{+0.01}$ $0.02_{-0.01}^{+0.01}$ $1.12_{-0.11}^{+0.14}$ $0.57_{-0.06}^{+0.06}$ $0.16_{-0.01}^{+0.01}$
@@ 5 Conclusions and discussion
Using the sample of 478 sGRBs with $P\geq 4.1$ $\mathrm{photons\,cm^{-2}\,s^{-1}}$ and $z\leq 3$, we investigated the sGRB LF and formation rate under various delay-time distribution models. We applied the maximum likelihood method to analyze the observed redshift and luminosity distributions of the sGRB sample, and used the MCMC technology to determine the best-fit parameters for each model. We found that all delay-time distribution models seem to account for the observed redshift and luminosity distributions similarly well. Nevertheless, based on the DIC model selection criterion, the Gaussian delay model is statistically preferred, with an estimated relative probability of $99.9\%$. Additionally, with the best-fit parameters presented in Table 2, consistent local formation rates of sGRBs are obtained for different delay-time distribution models, i.e., $R_{\mathrm{sGRB}}(0)=1.37_{-0.27}^{+0.30}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the Gaussian delay model, $1.63_{-0.27}^{+0.44}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the lognormal delay model, and $1.78_{-0.17}^{+0.12}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the power-law delay model. These results are well-aligned with those of other works (Guetta & Piran, 2005; Yonetoku et al., 2014; Ghirlanda et al., 2016; Paul, 2018).
@@ Acknowledgements
We thank the anonymous referee for his/her constructive comments, which greatly improved the clarity of this manuscript. We are also grateful to Dr. Shi-Jie Zheng, Dr. Guang-Xuan Lan, An Li, and Yi-Fang Liang for useful discussions. This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No. XDB0550400), the National Key R&D Program of China (2024YFA1611704), the National Natural Science Foundation of China (grant Nos. 12422307, 12373053, and 12321003), the Key Research Program of Frontier Sciences (grant No. ZDBS-LY-7014) of Chinese Academy of Sciences, and the Natural Science Foundation of Jiangsu Province (grant No. BK20221562).

### v3

@@ 4.1 GW detection rate simulation
We assume that the mass distribution of the NSs obeys a normal distribution with a mean of 1.33 $\mathrm{M_{\odot}}$ and a standard deviation of 0.09 $\mathrm{M_{\odot}}$ (Özel & Freire, 2016). The spins of the NSs are uniformly distributed between $0<\chi<0.05$, with the upper limit set by the maximum known NS spin (Zhu et al., 2018). The direction of the orbital angular momentum is uniformly distributed in space. We generate the GW waveforms using the Python package PyCBC [^3] https://pycbc.org/ with the IMPRPhenomPv2 approximant model (Schmidt et al., 2012; Hannam et al., 2014; Khan et al., 2016).
@@
$\mathrm{SNR_{det}^{2}}=4\operatorname{Re}\left(\int_{f_{\min}}^{f_{\max}}df\frac{\tilde{X}\tilde{X}^{*}}{S_{n}(f)}\right)\,$, (19)
@@
[…]
@@
Table 4: The sGRB and joint sGRB ++ GW detection rates for all combinations of high-energy GRB missions, GW networks, and delay-time distribution models. Columns 2-4 give the sGRB detection rates for the three GRB missions, independent of GW detection. Note that these rates are estimated for events with $d_{L}<1.6\,\mathrm{Gpc}$ (corresponding to $z<0.3$), which is the distance limit of this work; they should not be confused with the total sGRB detection rates (see the discussion at the end of section 4.2). The last six columns provide the joint sGRB ++ GW detection rates. For instance, the last cell of this table states that for the Gaussian delay model, EP/WXT ++ IGWN5 will detect $0.24_{-0.04}^{+0.03}$ events per year. All the results are presented at the $1\sigma$ confidence level.
@@
Gaussian $\mathrm{3.28_{-0.58}^{+0.48}}$ $2.69_{-0.40}^{+0.32}$ $0.78^{+0.09}_{-0.11}$ $24_{-0.03}^{+0.03} 0.0$$9_{-0.01}^{+0.01} 0.0$$02_{-0.01}^{+0.01}$ $1.18_{-0.21}^{+0.17}$ $0.69_{-0.12}^{+0.10}$ $0.24_{-0.04}^{+0.03}$
Lognormal $2.82_{-0.31}^{+0.22}$ $2.06_{-0.22}^{+0.16}$ $0.76_{-0.06}^{+0.08}$ $21_{-0.02}^{+0.03} 0.0$$8_{-0.01}^{+0.01} 0.0$$02_{-0.01}^{+0.01}$ $1.03_{-0.10}^{+0.19}$ $0.59_{-0.06}^{+0.10}$ $0.21_{-0.02}^{+0.01}$
Power law $3.19_{-0.35}^{+0.31}$ $1.99_{-0.25}^{+0.28}$ $0.57_{-0.07}^{+0.08}$ $21_{-0.02}^{+0.02} 0.0$$8_{-0.01}^{+0.01} 0.0$$02_{-0.01}^{+0.01}$ $1.12_{-0.11}^{+0.14}$ $0.57_{-0.06}^{+0.06}$ $0.16_{-0.01}^{+0.01}$
@@ 5 Conclusions and discussion
Using the sample of 478 sGRBs with $P\geq 4.1$ $\mathrm{photons\,cm^{-2}\,s^{-1}}$ and $z\leq 3$, we investigated the sGRB LF and formation rate under various delay-time distribution models. We applied the maximum likelihood method to analyze the observed redshift and luminosity distributions of the sGRB sample, and used the MCMC technology to determine the best-fit parameters for each model. We found that all delay-time distribution models seem to account for the observed redshift and luminosity distributions similarly well. Nevertheless, based on the DIC model selection criterion, the Gaussian delay model is statistically preferred, with an estimated relative probability of $99.9\%$. Additionally, with the best-fit parameters presented in Table 2, consistent local formation rates of sGRBs are obtained for different delay-time distribution models, i.e., $R_{\mathrm{sGRB}}(0)=1.37_{-0.27}^{+0.30}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the Gaussian delay model, $1.63_{-0.27}^{+0.44}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the lognormal delay model, and $1.78_{-0.17}^{+0.12}$ $\mathrm{Gpc^{-3}\,yr^{-1}}$ for the power-law delay model. These results are well-aligned with those of other works (Guetta & Piran, 2005; Yonetoku et al., 2014; Ghirlanda et al., 2016; Paul, 2018b).
@@
Finally, we note a caveat regarding our methodology. While we adopt the $E_{p}$-$L$ correlation for pseudo-redshift estimation throughout this work, significant systematic uncertainties affect this approach. Future refinement of redshift-estimation techniques or larger samples of sGRBs with spectroscopic redshifts will enable robust validation of our results.
@@ Acknowledgements
We thank the anonymous referee for their constructive comments, which greatly improved the clarity of this manuscript. We are also grateful to Dr. Shi-Jie Zheng, Dr. Guang-Xuan Lan, An Li, and Yi-Fang Liang for useful discussions. This work is partially supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No. XDB0550400), the National Key R&D Program of China (2024YFA1611704), the National Natural Science Foundation of China (grant Nos. 12422307, 12373053, and 12321003), and the Natural Science Foundation of Jiangsu Province (grant No. BK20221562).
@@
Appendix A Corner Plots of the MCMC inference
[…]
Figure 7: Same as Figure 6, but for the parameters of the Lognormal delay model.
Figure 8: Same as Figure 6, but for the parameters of the power-law delay model.